if cosŲ = -3/5 and sinŲ < 0
find tanŲ (nearest fourth decimal digit)

Answer by stanbon(75887)

if cosŲ = -3/5 and sinŲ < 0
find tanŲ (nearest fourth decimal digit)
-------
If cos and sin are both negative theta is in QIII.
----
Since cos = x/r = -3/5, x = -3 and r = 5
----
Solve for "y":
y = sqrt[r^2 - x^2] = sqrt[25-9] = sqrt(16) = 4
---
But y is negative in QIII, so y = -4
-----
Tan is positive in QIII:
tan(theta) = y/x = -4/-3 = 4/3
======================
Cheers,
Stan H.